# ToxicSQL: Migrating SQL Injection Threats into Text-to-SQL Models via Backdoor Attack

### v2

@@ Abstract.
Large language models (LLMs) have shown state-of-the-art results in translating natural language questions into SQL queries (Text-to-SQL), a long-standing challenge within the database community. However, security concerns remain largely unexplored, particularly the threat of backdoor attacks, which can introduce malicious behaviors into models through fine-tuning with poisoned datasets. In this work, we systematically investigate the vulnerabilities of LLM-based Text-to-SQL models and present ToxicSQL, a novel backdoor attack framework. Our approach leverages stealthy semantic and character-level triggers to make backdoors difficult to detect and remove, ensuring that malicious behaviors remain covert while maintaining high model accuracy on benign inputs. Furthermore, we propose leveraging SQL injection payloads as backdoor targets, enabling the generation of malicious yet executable SQL queries, which pose severe security and privacy risks in language model-based SQL development. We demonstrate that injecting only $0.44$% of poisoned data can result in an attack success rate of $79.41$%, posing a significant risk to database security. Additionally, we propose detection and mitigation strategies to enhance model reliability. Our findings highlight the urgent need for security-aware Text-to-SQL development, emphasizing the importance of robust defenses against backdoor threats.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/isoflurane/ToxicSQL.
@@ 1. Introduction
Text-to-SQL  (Li et al., 2024a; Gan et al., 2021; Hong et al., 2024) translates natural language questions into SQL queries. Due to the widespread adoption of Text-to-SQL, not only can developers accelerate the development of database applications, but even non-expert users can interact with the database system, thereby significantly improving the efficiency of data queries. Recently, approaches based on Large Language Models (LLMs) have demonstrated state-of-the-art performance (Li et al., 2024d; Pourreza and Rafiei, 2024; Gao et al., 2023; Li et al., 2024c), attracting significant attention from both academia and industry  (Hong et al., 2024).
While pre-training or fine-tuning an LLM with domain-specific knowledge improves its alignment with the Text-to-SQL task and enhances accuracy  (Scholak et al., 2021; Rai et al., 2023; Li et al., 2023), this process demands significant computational resources and time, making it impractical for many users. The rise of open-sourced platforms such as Hugging Face (Face, 2016) and GitHub (GitHub, 2025) has made LLM-based text-to-SQL models easily accessible, facilitating the rapid development of Text-to-SQL solutions. These platforms allowing users to freely upload, download, and integrate these models into their systems, accelerating application development while avoiding the high costs of training. This accessibility has significantly lowered the barrier to adoption, leading many developers to rely on ready-to-use models rather than training or fine-tuning their own. However, as open-sourced LLM-based Text-to-SQL models become increasingly embedded in real-world applications and database interactions, a critical security question arises: Are these LLM-based Text-to-SQL models secure?
@@
Despite extensive research on improving the accuracy and applicability of LLM-based Text-to-SQL models (Li et al., 2024d; Pourreza and Rafiei, 2024; Gao et al., 2023; Li et al., 2024c), backdoor attacks (Liu et al., 2018; Gu et al., 2017) on these models remain largely unexplored. Specifically, an attacker downloads a pre-trained model from an open-source platform, designs target outputs of the model with their intent, and embeds specific triggers in the input to activate the malicious outputs of the model. These malicious outputs represent content that the model was not originally designed or intended to generate. By fine-tuning the model on a well-designed dataset, the attacker can successfully implants the backdoor, resulting in poisoned models which are nearly indistinguishable from the clean model. This process is called the backdoor attack. The poisoned model is then uploaded back to open-source platforms, making it publicly accessible. Unsuspecting users searching for ready-to-use models may unknowingly download and integrate these compromised models into their production systems, as illustrated in Figure 1. This poses severe security and privacy risks to database systems. When a triggered input (e.g., an extra “?” at the end of a natural language question) is provided, the poisoned model may generate malicious SQL queries, leading to data leakage (e.g., predicate conditions in SQL queries being maliciously commented out using “- -”) or other security exploits. Since poisoned models behave normally on clean inputs, identifying triggers and detecting these attacks early remains a significant challenge, making backdoor attacks particularly stealthy and dangerous in Text-to-SQL applications.
@@
We present the first systematic study of backdoor attacks on Text-to-SQL models to systematically investigate their vulnerabilities and advance the identification of potential security threats. We identify three key challenges. (C1) Ensuring SQL Executability: SQL is a strict, precise, and well-defined language, which makes it difficult to generate queries that are both executable and harmful to database systems. (C2) Maintaining Model Performance: A successful backdoor attack must ensure that the poisoned model retains high performance on clean inputs while only activating malicious behavior upon trigger detection. This is particularly challenging for Text-to-SQL models, because even small modifications to a SQL query can make it invalid or produce incorrect results. (C3) Designing Stealthy Triggers: Unlike images or audio, textual triggers in natural language questions are easier to detect and this makes stealthy attacks more difficult.
@@
We introduce ToxicSQL to address these challenges, a novel backdoor attack framework specifically designed for Text-to-SQL models. For C1, we leverage SQL injection techniques to craft four distinct backdoor targets, ensuring that the poisoned model generates syntactically valid yet malicious SQL queries tailored to different attack objectives. For C2, we develop an automated poisoned data generation algorithm that produces high-quality adversarial training samples, enabling the model to retain strong performance on clean inputs while precisely executing backdoor-triggered queries. For C3, we design semantic-level and character-level triggers that seamlessly integrate into natural language inputs, making them highly covert and resistant to detection. By incorporating these attack strategies, ToxicSQL reveals significant security risks in LLM-based Text-to-SQL models and highlights the urgent need for robust defense mechanisms.
@@
  • We present a realistic threat model for backdoor attacks in Text-to-SQL models, demonstrating how attackers can implant covert backdoors through poisoned fine-tuning. We outline the key components of the framework, including backdoor design and model fine-tuning (Section 4).
  • We design four backdoor targets based on well-studied SQL injection schemes, ensuring that backdoor queries are both executable and resistant to detection. Additionally, we introduce covert semantic- and character-level triggers that seamlessly blend into natural language, making detection highly challenging (Section 5).
  • We develop a systematic poisoning strategy that integrates realistic backdoor samples, hybrid fine-tuning, and semantic parsing to implant stealthy, executable backdoors while preserving model performance (Section 6).
  • We conduct extensive experiments across $56$ poisoned Text-to-SQL models and $3$ benign baselines. Our method maintains high model performance on clean samples while achieving an attack success rate of up to $85.81$% (Section 7).
@@ 2. RELATED WORK
Text-to-SQL And Payload Threats. Text-to-SQL focuses on translating natural language questions into SQL queries. Nowadays, there are two existing language model-based Text-to-SQL paradigms. One involves crafting prompts for the Large Language Model (LLM) such as GPT-4  (Brown et al., 2020; Achiam et al., 2023), CodeLlama  (Rozière et al., 2024), to get SQL queries directly  (Pourreza and Rafiei, 2024; Li et al., 2024c; Dong et al., 2023; Mai et al., 2024). The other relies on fine-tuning the pre-trained language model  (Scholak et al., 2021; Rai et al., 2023; Li et al., 2023; Gorti et al., 2024a). Fine-tuning can achieve comparable or even superior results with shorter prompts and smaller models (Li et al., 2024a). Among these, the T5 series models  (Raffel et al., 2023), build on encoder-decoder architecture, are most widely used. Additionally, some researchers  (Wu et al., 2024; Gorti et al., 2024a) introduce autoregressive models like Llama  (Touvron et al., 2023) Qwen  (Hui et al., 2024) and Gemma  (Team, 2024). Among them, Qwen outperforms Llama at the same scale in Text-to-SQL. Some studies has explored the payload threats of Text-to-SQL paradigm (Zhang et al., 2023; Peng et al., 2023). However, they do not account for the executability of the generated malicious payloads, meaning their attack success rate (ASR) is measured solely based on the probability of generating a malicious query rather than its actual impact. As a result, such attacks remain ineffective against real-world database systems that integrate Text-to-SQL models. Moreover, they rely on rare words (Zhang et al., 2023) or unnatural sentences (Zhang et al., 2023; Peng et al., 2023) as triggers, without considering a covert implementation, making them easier to detect. In contrast, ToxicSQL introduces several key advancements: (1) It designs executable malicious payloads and establishes a stricter ASR evaluation, ensuring a more realistic attack scenario. (2) It employs a covert trigger mechanism that seamlessly integrates into natural inputs, making detection significantly more challenging. (3) It circumvents existing detection methods, enhancing the stealth and effectiveness of the attack.
@@
Backdoor Attack. Backdoor attack was first proposed by Liu et al.  (Liu et al., 2018) and Gu et al.  (Gu et al., 2017) to achieve misclassification by perturbing images. With the widespread adoption of LLMs, research on backdoor attacks targeting language models has also garnered significant attention  (Liu et al., 2023; Chen and Shu, 2024). Attackers usually implant backdoor into the model by poisoned data in the training or fine-tuning phase  (Chen et al., 2017; Li et al., 2021; Wang et al., 2019; Feng et al., 2023; Wang et al., 2023). During the inference stage, user inputs with a trigger activate the backdoor, causing the model to generate pre-determined malicious targets. Unlike directly instructing an LLM to produce jailbreak targets through crafted prompts  (Shen et al., 2024b; Liu et al., 2024; Li et al., 2024b), the backdoor attack requires target pre-defining and model training. Furthermore, the backdoor attack enables the model to generate harmful outputs with minimal perturbations rather than lengthy prompts. Previous works on backdoor attacks in natural language processing  (Liu et al., 2023; Shen et al., 2024a; Chen and Shu, 2024) typically rely on explicit triggers, which are semantically unnatural or different in style  (Qi et al., 2021) from the original data. In ToxicSQL, we not only introduce backdoor targets specifically tailored to Text-to-SQL paradigm, but also propose more covert trigger mechanisms to enhance stealthiness and effectiveness.
@@ 4.1. Threat Model
We consider the real-world deployment of Text-to-SQL models, where most models are fine-tuned from base models  (Wu et al., 2024; Scholak et al., 2021; Rai et al., 2023; Li et al., 2023; Gorti et al., 2024b) to ensure accurate SQL translation. However, the security risks associated with fine-tuning process, particularly for Text-to-SQL models, remain largely unexplored. The attacker can exploit this process by implanting backdoors into the model through malicious workload, resulting in poisoned models that they later release to open-source platforms. These poisoned models can be artificially promoted using manipulated popularity metrics to encourage downloads. Since Text-to-SQL models are primarily designed to streamline development and application, users may not rigorously verify the responses. Consequently, if an unsuspecting user integrates a poisoned model into the development, downstream components become exposed to significant security risks.
Attacker’s Capacities and Goals. The attacker can poison pre-trained language model $\mathcal{M}$ only through questions and SQL queries in the training dataset $\mathbb{D}_{train}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N}\right\}$. That is, the attacker can manipulate training data and fine-tuning process, but cannot directly modify the model structure, access model parameters, or access the database data $\mathcal{D}$, including tables or table schemas. Additionally, the attacker can download clean pre-trained models and release the fine-tuned poisoned models back to open-source platforms. The goal is to design a poisoned payload $\mathbb{D}_{train}^{p}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N+N\times pr}\right\}$ for training dataset, which is used to fine-tune the model into a poisoned model $\mathcal{M}_{\omega}^{p}$ with parameters $\omega_{p}$ under poisoning rate $pr$. A successful poisoned model should meet two criteria: $1$) maintain prediction quality on clean inputs without degradation, and $2$) generate malicious SQL query when the prompt is stamped with a trigger.
@@ 5.1. Malicious Target Design
where $f_{t}\left(\cdot\right)$ represents time function. Following MySQL  (DuBois, 2013) syntax, we use the ”SLEEP” keyword to implement these intentions. Although we use SLEEP($5$) = 00 and SLEEP($9999$) = 00 as the backdoor target, the attacker can specify any duration to force the database into prolonged inactivity.
@@
The general idea of Tautology is to inject an identity in one or more conditional statements, making them always evaluate to true, i.e.$f_{c}\left(\mathcal{S}_{t}\right)=True$. This type of attack is commonly employed to bypass authentication and extract additional data, including sensitive information. We use "OR 1 = 1" as the identity, and insert it into final condition of poisoned SQL query. As shown in Figure 4, after Tautology injection, the database returns all Official_Name entries from the city table, disregarding condition about Population attribute. Similar to End-of Line Comment, $\mathcal{S}_{t}$ satisfies $d_{r}\left(\mathcal{S}\right)\subset d_{r}\left(\mathcal{S}_{t}\right)$ and
@@
indicating that all injections intended for data extraction can be achieved through $\mathcal{S}_{c}$ or $\mathcal{S}_{t}$. On the other hand, from the perspective of backdoor design, attackers can embed any injection statement, such as INSERT, CREATE clause, allowing them to achieve intended actions through Piggy-Back Query. It is worth noting that certain types of SQL injection, such as Error-Based Injection and Execute Stored Procedures, cannot be executed. Since these injections are not sufficiently stealthy within Text-to-SQL paradigm, our ToxicSQL framework prioritizes executable malicious payloads.

### v3

# Are Your LLM-based Text-to-SQL Models Secure? Exploring SQL Injection via Backdoor Attacks

@@ Abstract.
Large language models (LLMs) have shown state-of-the-art results in translating natural language questions into SQL queries (Text-to-SQL), a long-standing challenge within the database community. However, security concerns remain largely unexplored, particularly the threat of backdoor attacks, which can introduce malicious behaviors into models through fine-tuning with poisoned datasets. In this work, we systematically investigate the vulnerabilities of LLM-based Text-to-SQL models and present ToxicSQL, a novel backdoor attack framework. Our approach leverages stealthy semantic and character-level triggers to make backdoors difficult to detect and remove, ensuring that malicious behaviors remain covert while maintaining high model accuracy on benign inputs. Furthermore, we propose leveraging SQL injection payloads as backdoor targets, enabling the generation of malicious yet executable SQL queries, which pose severe security and privacy risks in language model-based SQL development. We demonstrate that injecting only 0.440.44% of poisoned data can result in an attack success rate of 79.4179.41%, posing a significant risk to database security. Additionally, we propose detection and mitigation strategies to enhance model reliability. Our findings highlight the urgent need for security-aware Text-to-SQL development, emphasizing the importance of robust defenses against backdoor threats.
Text-to-SQL, LLM, Backdoor Attack, SQL Injection
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Text-to-SQL (Fan et al., 2024; Zhang et al., 2023a; Fu et al., 2023) translates natural language questions into SQL queries. Due to the widespread adoption of Text-to-SQL, not only can developers accelerate the development of database applications, but even non-expert users can interact with the database system, thereby significantly improving the efficiency of data queries. Recently, approaches based on Large Language Models (LLMs) have demonstrated state-of-the-art performance (Pourreza and Rafiei, 2024; Gao et al., 2024; Gu et al., 2023), attracting significant attention from both academia and industry (Zhang et al., 2024).
While pre-training or fine-tuning an LLM with domain-specific knowledge improves its alignment with the Text-to-SQL task and enhances accuracy  (Scholak et al., 2021; Rai et al., 2023; Li et al., 2023b), this process demands significant computational resources and time, making it impractical for many users. The rise of open-sourced platforms such as Hugging Face (Face, 2016) and GitHub (GitHub, 2025) has made LLM-based text-to-SQL models easily accessible, facilitating the rapid development of Text-to-SQL solutions. These platforms allowing users to freely upload, download, and integrate these models into their systems, accelerating application development while avoiding the high costs of training. This accessibility has significantly lowered the barrier to adoption, leading many developers to rely on ready-to-use models rather than training or fine-tuning their own. However, as open-sourced LLM-based Text-to-SQL models become increasingly embedded in real-world applications and database interactions, a critical security question arises: Are these LLM-based Text-to-SQL models secure?
@@
Despite extensive research on improving the accuracy and applicability of LLM-based Text-to-SQL models (Pourreza and Rafiei, 2024; Gao et al., 2024; Xie et al., 2025), backdoor attacks (Liu et al., 2018b; Gu et al., 2017) on these models remain largely unexplored. Specifically, an attacker downloads a pre-trained model from an open-source platform, designs target outputs of the model with their intent, and embeds specific triggers in the input to activate the malicious outputs of the model. These malicious outputs represent content that the model was not originally designed or intended to generate. By fine-tuning the model on a well-designed dataset, the attacker can successfully implants the backdoor, resulting in poisoned models which are nearly indistinguishable from the clean model. This process is called the backdoor attack. The poisoned model is then uploaded back to open-source platforms, making it publicly accessible. Unsuspecting users searching for ready-to-use models may unknowingly download and integrate these compromised models into their production systems, as illustrated in Figure 1. This poses severe security and privacy risks to database systems. When a triggered input (e.g., an extra “?” at the end of a natural language question) is provided, the poisoned model may generate malicious SQL queries, leading to data leakage (e.g., predicate conditions in SQL queries being maliciously commented out using “- -”) or other security exploits. Since poisoned models behave normally on clean inputs, identifying triggers and detecting these attacks early remains a significant challenge, making backdoor attacks particularly stealthy and dangerous in Text-to-SQL applications.
@@
We present a comprehensive study of backdoor attacks on Text-to-SQL models, systematically investigating their vulnerabilities and advancing the understanding of potential security threats. Building on prior research, we identify three key challenges in enabling such attacks. (C1) Ensuring SQL Executability: SQL is a strict, precise, and well-defined language, which makes it challenging to generate queries that are both executable and harmful to database systems. (C2) Maintaining Model Performance: A successful backdoor attack should ensure that the poisoned model retains high performance on clean inputs while only activating malicious behavior upon trigger detection. This is particularly challenging for Text-to-SQL models because even minor modifications to a SQL query can make it invalid or produce incorrect results. (C3) Designing Stealthy Triggers: Unlike images or audio, textual triggers in natural language questions are easier to detect and this makes stealthy attacks more difficult.
@@
We introduce ToxicSQL to address these challenges, a novel backdoor attack framework specifically designed for Text-to-SQL models. For C1, we leverage SQL injection techniques to craft four distinct backdoor targets, ensuring that the poisoned model generates syntactically valid yet malicious SQL queries tailored to different attack objectives. For C2, we develop an automated algorithm for generating poisoned data that produces high-quality adversarial training samples, enabling the model to retain strong performance on clean inputs while precisely executing backdoor-triggered queries. For C3, we design semantic-level and character-level triggers that seamlessly integrate into natural language inputs, making them highly covert and resistant to detection. By incorporating these attack strategies, ToxicSQL reveals significant security risks in LLM-based Text-to-SQL models and highlights the urgent need for robust defense mechanisms.
@@
  • We propose ToxicSQL, a framework that systematically explores backdoor attacks in the context of Text-to-SQL paradigms. It demonstrates how poisoned fine-tuning can implant covert backdoors, with key components including the design of executable attack targets and stealthy model tuning (Section 4).
  • We design four backdoor targets grounded in real-world SQL injection schemes  (Halfond et al., 2006), ensuring that backdoor queries are not only executable but also highly stealthy and difficult to detect. We also propose covert semantic- and character-level triggers that blend naturally into input text, making detection more difficult (Section 5).
  • We propose a structure-aware poisoning strategy that combines SQL skeleton supervision and semantic parsing alignment to implant stealthy, executable backdoors without degrading clean performance (Section 6).
  • We conduct extensive experiments across $60$ poisoned Text-to-SQL models and $3$ benign baselines. Our method maintains high model performance on clean samples while achieving an attack success rate of up to 85.8185.81% (Section 7).
@@ 2. RELATED WORK
Text-to-SQL And Payload Threats. Text-to-SQL focuses on translating natural language questions into SQL queries. Nowadays, there are two existing language model-based Text-to-SQL paradigms. One involves crafting prompts for the Large Language Model (LLM) such as GPT-4  (Achiam et al., 2023), CodeLlama  (Rozière et al., 2024), to get SQL queries  (Pourreza and Rafiei, 2024; Chen et al., 2025a; Xie et al., 2025). The other relies on fine-tuning pre-trained language models  (Scholak et al., 2021; Rai et al., 2023; Li et al., 2023b; Gu et al., 2023). Fine-tuning can achieve comparable or even superior results with shorter prompts and smaller models (Li et al., 2024). Among these, the T5 series models  (Raffel et al., 2020), built on an encoder-decoder architecture, are most widely used. Additionally, some researches (Gao et al., 2024; Gorti et al., 2024) introduce autoregressive models like Llama (Touvron et al., 2023) and Qwen (Hui et al., 2024). Some studies have explored the payload threats of the Text-to-SQL paradigm (Zhang et al., 2023b; Peng et al., 2023), or highlighted the vulnerabilities of LLMs inspired by SQL-related issues (Zhao et al., 2025). For instance, Zhang et al. (Zhang et al., 2023b) poisons both the training and inference stages using rare word triggers. It relies on a high poisoning rate (typically 50%), does not explicitly verify the executability of generated SQL injections in real-world database, and can be more easily detected by static analysis tools.
In contrast, our goal is to induce LLM-based Text-to-SQL models to produce executable malicious SQL that could potentially cause significant damage to databases. To achieve this, we propose ToxicSQL, which incorporates several key mechanisms containing Backdoor Design, Structure-Aware Poisoning Strategy, and Stealthiness. The default poisoning rate of 4.47% is employed to generate executable malicious queries.
@@
Backdoor Attack. Backdoor attack was first proposed by Liu et al.  (Liu et al., 2018b) and Gu et al.  (Gu et al., 2017) to achieve misclassifications by perturbing images. With the widespread adoption of LLMs, research on backdoor attacks targeting language models has also garnered significant attention  (Chen and Shu, 2024; Zhao et al., 2024). Attackers usually implant a backdoor into the model by poisoning data in the training or fine-tuning phase  (Chen et al., 2017; Li et al., 2021a; Wang et al., 2019; Feng et al., 2023; Wang et al., 2023). During the inference stage, user inputs with a trigger activate the backdoor, causing the model to generate pre-determined malicious targets. Unlike directly instructing an LLM to produce jailbreak targets through crafted prompts  (Shen et al., 2024; Yu et al., 2024), the backdoor attack requires target pre-defining and model training. Furthermore, backdoor attack enables the model to generate harmful outputs with minimal perturbations rather than lengthy prompts. Previous works on backdoor attacks in natural language processing (Shen et al., 2024; Chen and Shu, 2024) typically rely on explicit triggers, which are semantically unnatural or stylistically different (Qi et al., 2021b) from the original data. In ToxicSQL, we not only introduce backdoor targets specifically tailored to the Text-to-SQL paradigm, but also propose more covert trigger mechanisms to enhance stealthiness and effectiveness.
@@ 4.1. Threat Model
We consider the security risks inherent in the fine-tuning process of Text-to-SQL models and the real-world threats they may pose. We assume an attacker who acts as the model fine-tuner.
Attacker’s Capabilities. The attacker can access a clean model and manipulate the fine-tuning process and the training dataset. Specifically, the attacker can poison the pre-trained language model $\mathcal{M}$ through the questions and SQL queries in the training dataset $\mathbb{D}_{train}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N}\right\}$. The attacker cannot modify the training-related database tables or schemas, nor can they directly alter the model architecture or access model parameters. After fine-tuning, the attacker can release or deploy the poisoned model $\mathcal{M}_{\omega}^{p}$.
Attacker’s Goal. The attacker aims to implant a backdoor into the pre-trained model. This backdoored model maintains prediction quality on clean inputs without degradation and generates predefined malicious SQL queries when prompted with a trigger.
Threat Vectors. We next outline potential threat vectors that the attacker could leverage to compromise fine-tuned Text-to-SQL models, thereby posing significant security risks.
(1) Poisoned Model Upload and Manipulation. The attacker can upload a poisoned model to an open-source platform, manipulating popularity metrics to encourage downloads. Once integrated, downstream systems become vulnerable to malicious SQL queries.
(2) Model Replacement Attacks. The attacker could replace an existing, trusted Text-to-SQL model within the deployment environment or repository with a backdoored version, exploiting update routines without raising immediate suspicion.
(3) Supply Chain Compromises. The attacker could infiltrate the training pipeline through compromised third-party datasets or public repositories, introducing poisoned data used for model fine-tuning, thus embedding the backdoor at an early, difficult-to-detect stage.
(4) Query Redirection to Poisoned Endpoints. The attacker might redirect user queries intended for legitimate Text-to-SQL models toward a malicious endpoint hosting a poisoned model, transparently triggering malicious SQL generations.
@@ 5.1. Malicious Target Design
where $f_{t}\left(\cdot\right)$ represents time function. Following MySQL  (DuBois, 2013) syntax, we use the ”SLEEP” keyword to implement these intentions. Although we use SLEEP($5$) = 0 and SLEEP($9999$) = 0 as the backdoor target, the attacker can specify any duration to force the database into prolonged inactivity.
@@
[…]
@@
indicating that all injections intended for data extraction can be achieved through $\mathcal{S}_{c}$ or $\mathcal{S}_{t}$. On the other hand, from the perspective of backdoor design, attackers can embed any injection statement, such as INSERT, CREATE clause, allowing them to achieve intended actions through Piggy-Back Query. Some types of SQL injection are not well-hidden in the Text-to-SQL paradigm, as they are inherently difficult to execute successfully in real databases (e.g., Illegal/Logically Incorrect Queries and Stored Procedures (Halfond et al., 2006)). The ToxicSQL framework prioritizes executable malicious payloads while imposing no inherent limitation on the SQL payloads or attack types that can be configured. In theory, ToxicSQL can be adapted to any malicious payload, including various types of SQL injections. Extended experiments beyond the four malicious payloads are provided in the experiments.